The distance S in meters covered by a particle at time t seconds is given by

S=2t³-3t²-5t.Find

i) It's speed in the 5th second

ii) It's distance in the 3rd second

iii) The time the particle is momentarily at rest

iv) The acceleration at t=5s?


1
Expert's answer
2022-02-23T08:31:45-0500

Given:

"S=2t\u00b3-3t\u00b2-5t"

i) It's speed in the 5th second

"v=S'=6t^2-6t-5\\
v(5)=6*5^2-6*5-5=115\:\rm m/s"

ii) It's distance in the 3rd second

"S(3)=2*3^3-3*3^2-5*3=12\:\rm m"

iii) The time the particle is momentarily at rest

"v=6t^2-6t-5=0"

"t=1.54\:\rm s"

iv) The acceleration at t=5s?

"a=v'=12t-6\\
a(5)=12*5-6=54\:\rm m/s^2"